OpenStudy (anonymous):

solve the equation 4.9x+44=19.1

OpenStudy (juanpablojr):

Simplifying 4.9x + 44 = 19.1 Reorder the terms: 44 + 4.9x = 19.1 Solving 44 + 4.9x = 19.1 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-44' to each side of the equation. 44 + -44 + 4.9x = 19.1 + -44 Combine like terms: 44 + -44 = 0 0 + 4.9x = 19.1 + -44 4.9x = 19.1 + -44 Combine like terms: 19.1 + -44 = -24.9 4.9x = -24.9 Divide each side by '4.9'. x = -5.081632653 Simplifying x = -5.081632653